# FlowSteer: Interactive Agentic Workflow Orchestration via End-to-End Reinforcement Learning

### v2

@@ Abstract
In recent years, a variety of powerful agentic workflows have been applied to solve a wide range of human problems. However, existing workflow orchestration still faces key challenges, including high manual cost, reliance on specific operators/large language models (LLMs), and sparse reward signals. To address these challenges, we propose FlowSteer, an end-to-end reinforcement learning framework that takes a lightweight policy model as the agent and an executable canvas environment, automating workflow orchestration through multi-turn interaction. In this process, the policy model analyzes execution states and selects editing actions, while the canvas executes operators and returns feedback for iterative refinement. Moreover, FlowSteer provides a plug-and-play framework that supports diverse operator libraries and interchangeable LLM backends. To effectively train this interaction paradigm, we propose Canvas Workflow Relative Policy Optimization (CWRPO), which introduces diversity-constrained rewards with conditional release to stabilize learning and suppress shortcut behaviors. Experimental results on twelve datasets show that FlowSteer significantly outperforms baselines across various tasks. Our code is available at https://github.com/beita6969/FlowSteer.
@@
1The Chinese University of Hong Kong, Shenzhen  2Nanyang Technological University
@@
11footnotetext: Equal contribution.  Corresponding author.
@@ 1 Introduction
In recent years, a variety of powerful agentic systems have been applied to solve a wide range of human problems (Ruan et al., 2023; Shen et al., 2023; Hong et al., 2024), gradually moving beyond single-turn question answering (QA) toward executable end-to-end task completion. In this process, workflow orchestration has become a key bridge from task goals to reproducible execution: by organizing operators into an executable workflow graph, systems can complete complex tasks with improved controllability, debuggability, and reusability (Zeng et al., 2023; Qian et al., 2024), as shown in Figure 1. However, in practice, workflow construction still heavily relies on manual drag-and-drop and rule-based configuration (Hu et al., 2024), making it costly to transfer across new tasks, new operator libraries, new model backends, or different application domains.
@@
However, these methods still face several challenges: (i) High manual/heuristic dependence—rules and templates require continual maintenance and are often tightly coupled to specific scenarios, limiting reuse and generalization (Schick et al., 2023; Wang et al., 2024); (ii) Operator/backend lock-in (path lock-in)—existing approaches tend to rely on fixed operators or a single strong large language model (LLM) backend, making it difficult for a policy model to compose operators in a plug-and-play manner, so performance and robustness drop sharply when the operator library or backend environment changes (Ong et al., 2024; Chen et al., 2023); (iii) Sparse and unstable learning signals—training with only terminal correctness rewards can lead to shortcut behaviors (e.g., premature termination and oversimplified graphs) and reward hacking, and may further suffer from intra-group advantage collapse, making long-horizon credit assignment unstable (Yu et al., 2024; DeepSeek-AI, 2025).
@@
We evaluate FlowSteer on three task categories: QA, mathematical reasoning, and code generation. Experimental results demonstrate that FlowSteer enhances both generation quality and reasoning accuracy through its RL-driven multi-turn workflow interaction framework, surpassing baselines and existing methods, as shown in Figure 2. Beyond strengthening the reasoning capabilities of large-scale LLM backends, FlowSteer also improves the performance of smaller-scale LLMs, while reducing token consumption and interaction turns through more efficient orchestration strategies. Furthermore, FlowSteer can adapt across diverse task types without task-specific fine-tuning, demonstrating broad adaptability and strong practical potential.
@@ 4.2 Workflow Orchestration via Multi-Turn Interaction
Canvas-Feedback-Driven Workflow Interaction. Given action $a_{t}$ from Flow-Director, the Workflow Canvas executes the action and returns feedback $o_{t}^{\mathrm{exec}}$, forming an iterative closed-loop of “diagnose-edit-verify”. This interaction comprises three stages: (i) Action Execution and Feedback Generation. The canvas performs syntax parsing and constraint checking, then generates feedback $o_{t}^{\mathrm{exec}}$ including action success status, failure reasons, and repair suggestions:
@@
where $\mathcal{C}_{\mathrm{exec}}$ denotes the canvas feedback distribution. (ii) State Update and History Accumulation. The canvas updates the workflow graph (see Table 1 for update operations) and the state evolves via Eq. 5:
@@
(iii) Iterative Correction until Termination. Flow-Director continues decision-making based on $H_{t}$ until $\alpha_{t}=\texttt{finish}$ or $t=T_{\max}$, enabling error discovery and repair without manual rules. Trajectory Distribution and Optimization. Following Definition 2, a complete trajectory $\tau$ records the interaction from initialization to termination. The trajectory distribution is jointly determined by the policy and environment:
@@
You are building a workflow step by step to solve the problem. In each turn, output EXACTLY ONE XML action (add/delete/modify/set_prompt/finish or a structure action). The goal is to build a reliable workflow, not a single-shot answer. Keep your thinking brief and focus on choosing the next action, not solving the whole problem yourself. Let the operators do the computation. Proceed step by step with the following rules: <think> (brief reasoning for the current state and next action) </think> <action> (exactly one editing action with operator/target) </action> After the first step, in each interaction with the canvas, write: <think> (your reasoning based on the <feedback></feedback> from canvas) </think> <action> (new action to extend or refine the workflow) </action> Each <action> must build on what came before. Let the content of the workflow evolve naturally (for example: plan $\to$ programmer $\to$ verify $\to$ format). Continue producing think within <think></think> and action within <action></action> until the workflow is complete. Once the workflow is ready, write: <think> (final check that Format operator is added) </think> <action>finish</action> Task: {task}.
@@ 5.1 Experimental Setup
Datasets. We evaluate FlowSteer on six in-distribution (IID) benchmarks covering three task categories: (i) mathematical reasoning (GSM8K (Cobbe et al., 2021), MATH (Hendrycks et al., 2021b)), (ii) question answering (HotPotQA (Yang et al., 2018), SQuAD v2 (Rajpurkar et al., 2018)), and (iii) code generation (MBPP (Austin et al., 2021), HumanEval (Chen et al., 2021)). To assess generalization, we further test on six out-of-distribution (OOD) benchmarks: TriviaQA (Joshi et al., 2017), NaturalQuestions (Kwiatkowski et al., 2019), MathQA (Amini et al., 2019), AIME 2025 (AMC, 2025), APPS (Hendrycks et al., 2021a), and DS-1000 (Lai et al., 2023). More details of the datasets are provided in Appendix D.
@@ 5.3 Generalization Results (RQ2)
As shown in Table 4, FlowSteer maintains overall superiority over baseline methods across six OOD benchmarks, with more stable improvement trends in question answering and mathematical reasoning tasks. In comparison, directly using large-scale LLM backends typically exhibits stronger zero-shot capabilities but lacks transferable orchestration strategies, while search-based workflow methods are prone to efficiency limitations in OOD scenarios. These results demonstrate that FlowSteer achieves robust cross-task and cross-distribution improvements without task-specific fine-tuning of the backend LLM.
@@ 5.5 Ablation Study (RQ4)
As shown in Table 5, FlowSteer (Full) performs optimally across all datasets, highlighting the synergy of multi-turn interaction, executable feedback from Workflow Canvas, and end-to-end RL optimization. Removing any component significantly reduces performance, with RL most affected by complex reasoning tasks such as math and code generation. The canvas feedback has greater impact on QA tasks that require iterative error correction, showing task-dependent dependencies. In addition, FlowSteer (Full) exhibits more stable performance on the six OOD benchmarks, reflecting strong cross-task generalization ability. These results show that the effectiveness of FlowSteer stems from the close coupling of agent decision-making, canvas interaction, and RL optimization.
@@ 6 Conclusion
In this work, we propose FlowSteer, an end-to-end reinforcement learning framework for interactive workflow orchestration. By training a lightweight policy model to interact with an executable Workflow Canvas, FlowSteer learns transferable orchestration strategies from real execution feedback. We design diversity-constrained rewards with conditional release to jointly optimize workflow structure and task correctness. Experiments on twelve benchmarks demonstrate that FlowSteer significantly outperforms baselines across mathematical reasoning, QA, and code generation tasks.
@@ A.2 Operator Library $\mathcal{O}$
As shown in Table 7, the operator library $\mathcal{O}$ consists of 12 functional operators organized into six categories: planning (Plan, Decompose), solving (Programmer, Custom, AnswerGenerate), verification (Test, Review, Verify), revision (Revise), ensemble (ScEnsemble, Aggregate), and formatting (Format). The Programmer operator executes Python code in a sandboxed environment with scientific libraries under a 30-second timeout.
@@ E.2 Fine-Tuning Baselines
$\bullet$ SFT (Supervised Fine-Tuning) (Ouyang et al., 2022): A supervised fine-tuning baseline built on Qwen3-8B, trained with workflow annotation data to improve instruction following and DSL generation accuracy. Unlike Flow-Steer’s multi-turn interaction paradigm, SFT generates the complete workflow in a single turn without execution feedback. We use LoRA fine-tuning with rank 8 for parameter efficiency, evaluating how standard supervised adaptation enhances the raw backbone’s workflow construction capabilities.
$\bullet$ GRPO (Group Relative Policy Optimization) (Shao et al., 2024): Group Relative Policy Optimization is a reinforcement learning algorithm that normalizes rewards within sampled groups of trajectories. This reduces variance in policy updates, stabilizes training, and improves convergence efficiency compared to standard Proximal Policy Optimization (PPO). Unlike Flow-Steer’s multi-turn interaction with execution feedback, GRPO generates the entire workflow in one shot, limiting its ability to adapt based on intermediate results.
@@ E.4 Agent with Reinforcement Learning Methods
$\bullet$ Router-R1: A router-style architecture where a small policy model learns multi-round routing and aggregation of queries to different processing paths via reinforcement learning. The router iteratively selects from a predefined set of workflow templates with refinement across rounds. While this approach is computationally efficient, the routing mechanism has limited ability to adapt to intermediate execution results or refine workflows based on partial feedback, limiting its performance on complex multi-step reasoning tasks.

### v3

@@ Abstract
In recent years, a variety of powerful agentic workflows have been applied to solve a wide range of human problems. However, existing workflow orchestration still faces key challenges, including high manual cost, reliance on specific operators/LLMs, and sparse reward signals. To address these challenges, we propose FlowSteer, an end-to-end reinforcement learning framework that takes a lightweight policy model as the agent and an executable canvas environment, automating workflow orchestration through multi-turn interaction. In this process, the policy model analyzes execution states and selects editing actions, while the canvas executes operators and returns feedback for iterative refinement. Moreover, FlowSteer provides a plug-and-play framework that supports diverse operator libraries and interchangeable LLM backends. To effectively train this interaction paradigm, we propose Canvas Workflow Relative Policy Optimization (CWRPO), which introduces diversity-constrained rewards with conditional release to stabilize learning and suppress shortcut behaviors. Experimental results on twelve datasets show that FlowSteer significantly outperforms baselines across various tasks. Our code is available at https://github.com/beita6969/FlowSteer.
@@
1The Chinese University of Hong Kong, Shenzhen
2Nanyang Technological University
@@
footnotetext: Equal contribution.  Corresponding author.
@@ 1 Introduction
In recent years, a variety of powerful agentic systems have been applied to solve a wide range of human problems (Ruan et al., 2023; Shen et al., 2023; Hong et al., 2024), gradually moving beyond single-turn question answering (QA) toward executable end-to-end task completion.
In this process, workflow orchestration has become a key bridge from task goals to reproducible execution: by organizing operators into an executable workflow graph, systems can complete complex tasks with improved controllability, debuggability, and reusability (Zeng et al., 2023; Qian et al., 2024), as shown in Figure 1. However, in practice, workflow construction still heavily relies on manual drag-and-drop and rule-based configuration (Hu et al., 2024), making it costly to transfer across new tasks, new operator libraries, new model backends, or different application domains.
@@
However, these methods still face several challenges:
(i) High manual/heuristic dependence—rules and templates require continual maintenance and are often tightly coupled to specific scenarios, limiting reuse and generalization (Schick et al., 2023; Wang et al., 2024);
(ii) Operator/backend lock-in (path lock-in)—existing approaches tend to rely on fixed operators or a single strong large language model (LLM) backend, making it difficult for a policy model to compose operators in a plug-and-play manner, so performance and robustness drop sharply when the operator library or backend environment changes (Ong et al., 2024; Chen et al., 2023);
(iii) Sparse and unstable learning signals—training with only terminal correctness rewards can lead to shortcut behaviors (e.g., premature termination and oversimplified graphs) and reward hacking, and may further suffer from intra-group advantage collapse, making long-horizon credit assignment unstable (Yu et al., 2024; DeepSeek-AI, 2025).
@@
We evaluate FlowSteer on three task categories: QA, mathematical reasoning, and code generation. Experimental results demonstrate that FlowSteer enhances both generation quality and reasoning accuracy through its RL-driven multi-turn workflow interaction framework, surpassing baselines and existing methods, as shown in Figure 2.
Beyond strengthening the reasoning capabilities of large-scale LLM backends, FlowSteer also improves the performance of smaller-scale LLMs, while reducing token consumption and interaction turns through more efficient orchestration strategies. Furthermore, FlowSteer can adapt across diverse task types without task-specific fine-tuning, demonstrating broad adaptability and strong practical potential.
@@ 4.2 Workflow Orchestration via Multi-Turn Interaction
Canvas-Feedback-Driven Workflow Interaction. Given action $a_{t}$ from Flow-Director, the Workflow Canvas executes it and returns feedback $o_{t}^{\mathrm{exec}}$, forming a closed-loop “diagnose-edit-verify”. This interaction has three stages: (i) Action Execution and Feedback Generation. The canvas performs syntax parsing and constraint checking, then provides $o_{t}^{\mathrm{exec}}$ with success status, failure reasons, and repair suggestions:
@@
where $\mathcal{C}_{\mathrm{exec}}$ denotes the canvas feedback distribution. (ii) State Update and History Accumulation. The canvas updates the workflow graph (Table 1) and the state evolves via Eq. 5:
@@
(iii) Iterative Correction until Termination. Flow-Director continues decision-making based on $H_{t}$ until $\alpha_{t}=finish$ or $t=T_{\max}$, enabling error discovery and repair without manual rules. Trajectory Distribution and Optimization. Following Definition 2, a complete trajectory $\tau$ records the interaction from initialization to termination. The trajectory distribution is jointly determined by the policy and environment:
@@
You are building a workflow step by step to solve the problem. In each turn, output EXACTLY ONE XML action (add/delete/modify/set_prompt/finish or a structure action). The goal is to build a reliable workflow, not a single-shot answer. Keep your thinking brief and focus on choosing the next action, not solving the whole problem yourself. Let the operators do the computation. Proceed step by step with the following rules: <think> (brief reasoning for the current state and next action) </think> <action> (exactly one editing action with operator/target) </action> After the first step, in each interaction with the canvas, write: <think> (your reasoning based on the <feedback></feedback> from canvas) </think> <action> (new action to extend or refine the workflow) </action> Each <action> must build on what came before. Let the content of the workflow evolve naturally (for example: plan $\to$ solve $\to$ verify $\to$ format). Continue producing think within <think></think> and action within <action></action> until the workflow is complete. Once the workflow is ready, write: <think> (final check that Format operator is added) </think> <action>finish</action> Task: {task}.
@@ 5.1 Experimental Setup
Datasets. We evaluate FlowSteer on diverse in-distribution (IID) and out-of-distribution (OOD) benchmarks, compare against strong baselines, and report standard task-specific metrics.
We use six IID benchmarks covering three task categories: (i) mathematical reasoning (GSM8K (Cobbe et al., 2021), MATH (Hendrycks et al., 2021b)), (ii) question answering (HotPotQA (Yang et al., 2018), SQuAD v2 (Rajpurkar et al., 2018)), and (iii) code generation (MBPP (Austin et al., 2021), HumanEval (Chen et al., 2021)). To assess generalization, we also test on six OOD benchmarks: TriviaQA (Joshi et al., 2017), NaturalQuestions (Kwiatkowski et al., 2019), MathQA (Amini et al., 2019), AIME 2025 (AMC, 2025), APPS (Hendrycks et al., 2021a), and DS-1000 (Lai et al., 2023). More details in Appendix D.
@@ 5.3 Generalization Results (RQ2)
As shown in Table 4, FlowSteer maintains superiority over baselines on six OOD benchmarks, with stable improvements in question answering and mathematical reasoning. In contrast, large-scale LLM backends show strong zero-shot capability but lack transferable orchestration, while search-based workflows are inefficient in OOD settings. These results show that FlowSteer achieves robust cross-task and cross-distribution gains without task-specific fine-tuning of the backend LLM.
@@ 5.5 Ablation Study (RQ4)
As shown in Table 5, FlowSteer (Full) performs optimally across all datasets, highlighting the synergy of multi-turn interaction, executable feedback from Workflow Canvas, and end-to-end RL optimization. Removing any component significantly reduces performance, with RL most affected by complex reasoning tasks such as math and code generation. The canvas feedback has greater impact on QA tasks that require iterative error correction, showing task-dependent dependencies.
In addition, FlowSteer (Full) exhibits more stable performance on the six OOD benchmarks, reflecting strong cross-task generalization ability. These results show that the effectiveness of FlowSteer stems from the close coupling of agent decision-making, canvas interaction, and RL optimization.
@@ 6 Conclusion
In this work, we propose FlowSteer, an end-to-end reinforcement learning framework for interactive workflow orchestration. By training a lightweight policy model to interact with an executable Workflow Canvas, FlowSteer learns transferable orchestration strategies from real execution feedback.
We design diversity-constrained rewards with conditional release to jointly optimize workflow structure and task correctness. Experiments on twelve benchmarks demonstrate that FlowSteer significantly outperforms baselines across mathematical reasoning, QA, and code generation tasks.
@@ A.2 Operator Library $\mathcal{O}$
As shown in Table 7, the operator library $\mathcal{O}$ includes 12 operators in six categories: planning (Plan, Decompose), solving (Programmer, Custom, AnswerGenerate), verification (Test, Review, Verify), revision (Revise), ensemble (ScEnsemble, Aggregate), and formatting (Format). The Programmer operator executes Python code in a sandboxed environment with scientific libraries under a 30-second timeout.
@@ E.2 Fine-Tuning Baselines
$\bullet$ Supervised Fine-Tuning (Ouyang et al., 2022): An SFT baseline built on Qwen3-8B, trained with workflow annotation data to improve instruction following and DSL generation accuracy. Unlike Flow-Steer’s multi-turn interaction paradigm, SFT generates the complete workflow in a single turn without execution feedback. We use LoRA fine-tuning with rank 8 for parameter efficiency, evaluating how standard supervised adaptation enhances the raw backbone’s workflow construction capabilities.
$\bullet$ Group Relative Policy Optimization (Shao et al., 2024): GRPO is a reinforcement learning algorithm that normalizes rewards within sampled groups of trajectories. This reduces variance in policy updates, stabilizes training, and improves convergence efficiency compared to standard Proximal Policy Optimization (PPO). Unlike Flow-Steer’s multi-turn interaction with execution feedback, GRPO generates the entire workflow in one shot, limiting its ability to adapt based on intermediate results.
@@ E.4 Agent with Reinforcement Learning Methods
$\bullet$ Router-R1: A router-style architecture where a small policy model learns multi-round routing and aggregation of queries to different processing paths via reinforcement learning. The router iteratively selects from a predefined set of workflow templates with refinement across rounds. While this approach is computationally efficient, the routing mechanism has limited ability to adapt to intermediate execution results or refine workflows on partial feedback, limiting its performance on complex multi-step reasoning.